# Impacts of Data Splitting Strategies on Parameterized Link Prediction Algorithms

### v1

@@ Acknowledgments
More than a decade ago, one of the authors of this paper (Tao Zhou) had an argument with Dr. Junming Huang regarding whether the two-set split or the three-set split should be used in link prediction. Caused by naivety and a lack of experience in artificial intelligence, Tao Zhou believed that the two-set split was feasible, which was also the mainstream method at that time. They co-authored a paper titled "Predicting missing links via significant paths" [EPL 106 (2014) 18008] using the two-set splitting strategy. Notably, the parameterized algorithm used as a benchmark in that Letter was the local path index, whose results were surprisingly close in the two data splits (as shown in this paper). Due to their differing viewpoints, Dr. Junming Huang later withdrew his name as a co-author, as indicated in the erratum [Erratum: Predicting missing links via significant paths, EPL 108 (2014) 49901]. The results of this paper strongly demonstrate that Dr. Junming Huang was correct! This work was partially supported by the National Natural Science Foundation of China (Nos. T2293771 and 42361144718) and the STI 2030–Major Projects (No. 2024ZD0 523903).

### v2

@@ 4 Results
Figure 2: Robustness of the loss ratio $\mathcal{L}$ with respect to random data partitioning. The figure summarizes the distribution of the standard deviation $\mathrm{Std}(\mathcal{L})$ for each algorithm across all settings. For each setting, $\mathrm{Std}(\mathcal{L})$ is computed from multiple independent repetitions of the data splitting procedure, thereby quantifying the variability induced by partition randomness. Lower values indicate stronger statistical robustness. Boxplots report the median and interquartile range, while individual scatter points correspond to specific combinations of network and $\rho$.
@@
We further examine the statistical robustness of the loss ratio $\mathcal{L}$ with respect to the randomness of data partitioning. Specifically, for each setting (i.e., algorithm + network + $\rho$), we repeat the data splitting procedure multiple times and compute the resulting standard deviation $\mathrm{Std}(\mathcal{L})$. Figure 2 shows the distribution of $\mathrm{Std}(\mathcal{L})$ across all settings. We observe that the variability of $\mathcal{L}$ is consistently low for the majority of algorithms. In particular, classical heuristic and random-walk-based methods exhibit near-zero standard deviations across almost all settings, indicating strong robustness to random partitioning. Learning-based methods display moderately higher variability, yet the fluctuations are still well bounded and do not affect the qualitative conclusions. Overall, these results demonstrate that the observed performance gap between two-set and three-set splitting strategies is statistically stable, rather than an artifact arising from random partition.
@@ Acknowledgments
More than a decade ago, one of the authors of this paper (Tao Zhou) had an argument with Dr. Junming Huang regarding whether the two-set split or the three-set split should be used in link prediction. Caused by naivety and a lack of experience in artificial intelligence, Tao Zhou believed that the two-set split was feasible, which was also the mainstream method at that time. They co-authored a paper titled ”Predicting missing links via significant paths” [EPL 106 (2014) 18008] using the two-set splitting strategy. Notably, the parameterized algorithm used as a benchmark in that Letter was the local path index, whose results were surprisingly close in the two data splits (as shown in this paper). Due to their differing viewpoints, Dr. Junming Huang later withdrew his name as a co-author, as indicated in the erratum [Erratum: Predicting missing links via significant paths, EPL 108 (2014) 49901]. The results of this paper strongly demonstrate that Dr. Junming Huang was correct! This work was partially supported by the National Natural Science Foundation of China (Nos. T2293771 and 42361144718) and the STI 2030–Major Projects (No. 2024ZD0523903). Y.L. acknowledges financial support from the University of Warwick and the China Scholarship Council (CSC) under the Scholarship Fund (File No. 202508060219).